Can you help me with this exercise?
cota-tana-2tan2a=4cot4a

Grade:10

3 Answers

yash lohar
7 years ago
To solve this problem simply you should use the formula of cot(a+b)=1/tan(a+b) So use the formula of cot(a+b) to expand cot(2a+2a) and then separate ly divide the numerator by the denominator tan2a and then again to eliminate cot2a so formed use the relation cot(a+a) and follow the same method as done above
sachin
6 years ago
To solve this problem simply you should use the formula of cot(a+b)=1/tan(a+b) So use the formula of cot(a+b) to expand cot(2a+2a) and then separate ly divide the numerator by the denominator tan2a and then again to eliminate cot2a so formed use the relation cot(a+a) and follow the same method as done above,
sachin
6 years ago
To solve this problem simply you should use the formula of cot(a+b)=1/tan(a+b) So use the formula of cot(a+b) to expand cot(2a+2a) and then separate ly divide the numerator by the denominator tan2a and then again to eliminate cot2a so formed use the relation cot(a+a) and follow the same method as done above,,,,